Suppose the firm’s total cost and marginal cost functions are given by TC=54+Q+2Q^3 and MC=1+4Q^2, respectively. What is the output level that minimizes average total cost?

A. 2

B. 3

C. 6

D. 8
Answer #1

Average total cost= 54/Q+1+2Q^2.

the average cost is minimum where it is equal to marginal cost.

1+4Q^2 = 54/Q +1+2Q^2

2Q^2 = 54/Q

Q^3 = 27

Q= 3.

Hence the correct option is

B. 3.